OpenStudy (anonymous):

y^2-x^2=1, y=2; about the y-axis. Find the volume of the resulting solid by any method.

OpenStudy (anonymous):

Is there a lower boundary? Are you only considering the top half?

OpenStudy (anonymous):

Only the area where the two lines intersect.

OpenStudy (anonymous):

Gotcha. Ok, I think you can do a disk method here - sum of many cylinders. Volume of cylinder = pir*r^2*h, r=x=sqrt(y^2-1), and h=dy. Integrate between y=1 and y=2. That should do it.

OpenStudy (anonymous):

*(pir should be pi)

OpenStudy (anonymous):

Cool, I'm going to compute that and see if I get a plausible answer. Thank You

OpenStudy (anonymous):

Ok, let me know what you get so I can check my answer.

OpenStudy (anonymous):

I get 4pi/3 or 4.19

OpenStudy (anonymous):

Same here. You can do a reasonableness check by comparing it to a similarly sized cone and use the volume formula for a cone.